Constants SubmitP The electric field at a distance of 0.154 m from the surface of a solid insulating sphere with radius 0.364 m is 1750 N/C X Incorrect; Try Again; One attempt remaining You may want to review (Pages 738-742). For related problem-solving tips and strategies, you may want to view a Video Tutor Solution of Field of a uniformly charged sphere. Part B Calculate the electric field inside the sphere at a distance of 0.241 Irfrom the center. 0图 N/C Submit Request Answer Next > Provide Feedback Previous
Answer #1

Solution)

Given, Radius of the sphere R = 0.364 m

Distance of the point from the surface of the sphere x = 0.154 m

Distance of the point from center of the sphere r = R + x

= 0.518 m

Electric field , E = 1750 N / C

We know,

E = Kq / r ^2

So, q = Er^ 2 / K

Now, on substituting values we get

q = 5.217 x 10 ^-8 C

Part b)

We know,

Distance r ' = 0.241 m

Where, R=0.364 m

Electric field E = Kqr ' / R ^3

On substituting values, we get

E= 2346.26 N/C (Ans)

========

Good luck!:)